### v1

@@ 1 Introduction
The next-generation Stage IV survey telescopes are expected to provide such capabilities and means, such as $Euclid$ [2], Nancy Grace Roman Space Telescope (RST) [^1] https://roman.gsfc.nasa.gov/, Vera C. Rubin Observatory [3], and the China space station telescope (CSST) [4, 5, 6]. These telescope have large field of view and high spatial resolution and sensitivity, which will explore huge volumes of space covering vast redshift range and observe billions of galaxies. They can measure the property of the components of the Universe, formation and evolution of cosmic large-scale structure (LSS), and expansion history of the Universe with extremely high accuracy.
@@
Magnitude limitb (multi/single) $28.0/25.3$ $28.0/25.7$ $28.7/26.4$ $28.4/26.1$ $28.2/25.9$ $27.7/25.4$ $27.1/24.4$
@@ 2.1 Weak gravitational lensing
In [14], by dividing the redshift range from $z=0$ to 3 into four tomographic bins, the mock data of the auto and cross convergence power spectra between different bins are derived. The galaxy photo-$z$ distributions in the four bins are shown in Figure 2. The galaxy photometric redshift can be well estimated by the SED template fitting methods or mechine learning with the accuracy better than $\sigma_{z0}\simeq 0.03$ [15, 16, 17, 18, 19, 20, 21]. Considering the cosmological parameters, such as the reduced Hubble constant $h$, matter density parameter $\Omega_{\rm m}$, baryon density parameter $\Omega_{\rm b}$, the amplitude parameter of matter fluctuation on a scale of 8 $h^{-1}$pc $\sigma_{8}$, spectral index $n_{\rm s}$, and the equation of stat of dark energy $w$, the intrinsic alignment, i.e. $A_{\rm IA}$ and $\eta_{\rm IA}$, and photo-$z$ shift parameter $\Delta z^{i}$, photo-$z$ stretch parameter $\sigma_{z}^{i}$, and shear calibration parameter $m_{i}$ in each tomographic bin, they performed the Markov Chain Monte Carlo (MCMC) to jointly constrain the model parameters.
@@ 3.1 Type Ia supernovae
Besides the CSST-UDF survey, the CSST wide-field survey also can be used to detect a large number of SNe in ten years, although it is not dedicated to time-domain studies. As shown in [45], about 5 million SNe of various types can be detected by CSST for follow-up observations by time-domain telescopes. This will greatly promote future SN studies, and can be helpful to measure the expansion history of the Universe with extreme accuracy.
@@ 4 Discussion and conclusions
In this work, we review the main cosmological probes in the CSST photometric and spectroscopic surveys and the constraint power on dark energy, dark matter, and modified gravity theories. We find that, by implementing the weak gravitational lensing, 2D and 3D galaxy clustering, galaxy cluster abundance, cosmic void, SN Ia, and BAO measurements, CSST can perform joint constraints on cosmological parameters and obtain extremely precise results by effectively breaking the parameter degeneracies.
CSST will measure the equation of state of dark energy with an accuracy better than 5% and even close to $\sim$1%, which can determine if dark energy is the cosmological constant or has dynamical evolution as some scalar field or other models. The property of dark matter, e.g. cold or warm, can be clearly distinguished, and the dark matter particle mass (e.g. axion) will be precisely measured, which will provide important reference for detecting dark matter particles in the laboratory. Besides, the theory of gravity can be tested in high precision, and modified gravity models, e.g. Brans-Dicke theory and $f(R)$ theory, could be stringently constrained.
In addition to the probes discussed above, CSST also can perform other cosmological probes, e.g. the strong gravitational lensing [49, 50], 2D BAO [51], AP effect [52], cosmic optical background [53], etc. Moreover, CSST also has strong synergy with other telescopes, such as other Stage IV survey telescopes like $Euclid$ [54], future CMB experiments [55], time-domain telescopes for detecting transient sources [45], the radio telescopes like FAST [56] and MeerKAT [57] for extracting 21cm signal, and future gravitational wave detectors [58]. All these indicate that CSST can greatly promote the relevant cosmological studies in the future, and would provide new physics and opportunities for cosmology.

### v2

@@ 1 Introduction
The next-generation Stage IV survey telescopes are expected to provide such capabilities and means, such as $Euclid$ [2], Nancy Grace Roman Space Telescope (RST) [^1] https://roman.gsfc.nasa.gov/, Vera C. Rubin Observatory [3], and the China space station telescope (CSST) [4, 5, 6]. These telescopes have large field of view and high spatial resolution and sensitivity, which will explore huge volumes of space covering vast redshift range and observe billions of galaxies. They can measure the property of the components of the Universe, formation and evolution of cosmic large-scale structure (LSS), and expansion history of the Universe with extremely high accuracy.
@@
Mag. limitb (multi/single expo.) $28.0/25.3$ $28.0/25.7$ $28.7/26.4$ $28.4/26.1$ $28.2/25.9$ $27.7/25.4$ $27.1/24.4$
@@ 2.1 Weak gravitational lensing
In [14], by dividing the redshift range from $z=0$ to 3 into four tomographic bins, the mock data of the auto and cross convergence power spectra between different bins were derived. The galaxy photo-$z$ distributions in the four bins are shown in Figure 2. The galaxy photometric redshift can be well estimated by the SED template fitting methods or mechine learning with the accuracy better than $\sigma_{z0}\simeq 0.03$ [15, 16, 17, 18, 19, 20, 21]. Considering the cosmological parameters, such as the reduced Hubble constant $h$, matter density parameter $\Omega_{\rm m}$, baryon density parameter $\Omega_{\rm b}$, the amplitude parameter of matter fluctuation on a scale of 8 $h^{-1}$pc $\sigma_{8}$, spectral index $n_{\rm s}$, and the equation of stat of dark energy $w$, the intrinsic alignment, i.e. $A_{\rm IA}$ and $\eta_{\rm IA}$, and photo-$z$ shift parameter $\Delta z^{i}$, photo-$z$ stretch parameter $\sigma_{z}^{i}$, and shear calibration parameter $m_{i}$ in each tomographic bin, they performed the Markov Chain Monte Carlo (MCMC) to jointly constrain the model parameters.
@@ 3.1 Type Ia supernovae
Besides the CSST-UDF survey, the CSST wide-field survey also can be used to detect a large number of SNe in ten years, although it is not dedicated to time-domain studies. As shown in [45], about 5 million SNe of various types can be detected by CSST for follow-up observations by time-domain telescopes. This will greatly promote future SN studies, which can be helpful to measure the Hubble constant $H_{0}$ and paramter $H(z)$, and probe the expansion history of the Universe with extreme accuracy.
@@ 4 Discussion and conclusions
In this work, we review the main cosmological probes in the CSST photometric and spectroscopic surveys and the constraint power on dark energy, dark matter, and modified gravity theories. CSST can precisely measure matter distribution of the Universe by probing weak gravitational lensing, 2D and 3D galaxy clustering, galaxy cluster abundance and cosmic void, and detect the cosmic expansion history by SN Ia, and BAO observations. By implementing these measurements, CSST will perform joint constraints on cosmological parameters and obtain extremely precise results by effectively breaking the parameter degeneracies.
We find that CSST can measure the equation of state of dark energy with an accuracy better than 5% and even close to $\sim$1%, which can determine if dark energy is the cosmological constant or has dynamical evolution as some scalar field or other models. The property of dark matter, e.g. cold or warm, can be clearly distinguished, and the dark matter particle mass (e.g. axion) will be precisely measured, which will provide important reference for detecting dark matter particles in the laboratory. Besides, the theory of gravity can be tested in high precision, and modified gravity models, e.g. Brans-Dicke theory and $f(R)$ theory, could be stringently constrained.
On the other side, it is worth noting that combining probes is not a trivial task, since different probes may suffer different kinds of systematics, and their results may have tension. This brings a challenge to the joint cosmological constrints of CSST and other Stage IV surveys. However, by adopting more advanced techniques of the data processing and analysis, e.g. mechine learning, we expect that this issue can be well solved. Besides, many parameters in the analyses mentioned above are adjustable (e.g., the number of redshift bins and the range of each bin, the scale range of galaxy clustering analysis, the definition of cosmic voids, etc.). Adjusting these options might further enhance the constraint power, and improve the fitting accuracy.
In addition to the probes discussed above, CSST also can perform other cosmological probes, e.g. the strong gravitational lensing [49, 51], 2D BAO [52], AP effect [53], cosmic optical background [54], galaxy-ellipticity correlation [50], etc. Moreover, CSST also has strong synergy with other telescopes, such as other Stage IV survey telescopes like $Euclid$ [55], future CMB experiments [56], time-domain telescopes for detecting transient sources [45], the radio telescopes like FAST [57] and MeerKAT [58] for extracting 21cm signal, and future gravitational wave detectors [59]. All these indicate that CSST can greatly promote the relevant cosmological studies in the future, and new physics and opportunities related to new discoveries and theories about dark energy, dark matter, and modified gravity are expected to emerge in abundance.